Is this statement true about enzyme?
“enzyme-substrate complexes with affinity, must have a low Vmax”
Answer #1

The enzyme affinity for the substrate is measured as:

  • KM = KS + (K2/K1)
  • KS = (K-1/K1) = [E] [S]/ [ES] = Enzymes Substrate dissociation constant

Michaelis-Menten equation may be derived as:

V0 = Vmax [S]/ KM+[S]

  • Vmax = At maximum substrate concentration or saturation, the maximum velocity obtained.
  • KM = Michaelis constant: It is described as the concentration of substrate concentration when the velocity is half or 50% of maximum velocity. OR, Vmax/2
  • [S] = concentration of the substrate.

At lower KM the binding affinity increases. Or Km is inversely proportional to the affinity of substrate to form enzyme-substrate complex.